OpenStudy (anonymous):

integrate dx/(x+2)square root x^2+4x+8

OpenStudy (hba):

\[\int\limits_{}^{}\frac{ dx }{ x+2 }\sqrt{x^2+4x+8}\] Right ?

OpenStudy (anonymous):

square root at the bottom

OpenStudy (anonymous):

i"ll try to do partial fraction but i got o for a

OpenStudy (hba):

Use substitution :)

OpenStudy (anonymous):

this is my working.,but it does not make sense., \[\int\limits_{}^{} \frac{ dx}{ (x+2)(\sqrt{(x+2)^{2}} }\]

OpenStudy (anonymous):

is it correct if i just cancel out the squaare root

hartnn (hartnn):

use (x+2)/2 = tan y

OpenStudy (hba):

\[\int\limits_{}^{}\frac{ 1 }{ x+2\sqrt{x^2+4x+8} }=\int\limits_{}^{}\frac{ 1 }{ x+2\sqrt{(x+2)^2 +4} }\] Substiute u=x+2 du=dx

OpenStudy (hba):

Sorry i missed the dx :/

hartnn (hartnn):

(x+2)^2 is not equal to x^2+4x+8

OpenStudy (hba):

I wrote (x+2)^2+4=x^2+4x+8

hartnn (hartnn):

i said that to gee..

OpenStudy (hba):

\[\int\limits_{}^{}\frac{ 1 }{ u+\sqrt{u^2+4 } }du\] Now substitute u=2tan(y) and du=2sec^2 ydy

OpenStudy (anonymous):

@hartnn ouh ok.,i did not noticce it is polar

OpenStudy (shubhamsrg):

@hba its multiplication in the denominator there..

hartnn (hartnn):

\(\int\limits_{}^{} \frac{ dx}{ 2[(x+2)/2](\sqrt{((x+2)/2)^{2}+1} }\) now try tan y = (x+2)/2

hartnn (hartnn):

you should get something of the form, c cosec y

OpenStudy (shubhamsrg):

1/usqrt(u^2 +4) =1/(u^2) sqrt( 1 + 4/u^2) let 1/u = t no need for trig substitution if your not comfortable with it..

OpenStudy (shubhamsrg):

it then reduces to standard form..

OpenStudy (kainui):

Useful: \[\csc \theta *\frac{ \csc \theta + \cot \theta }{ \csc \theta + \cot \theta }=\frac{ \csc^2 \theta + \cot \theta \csc \theta }{ \cot \theta + \csc \theta }\]

OpenStudy (anonymous):

@hba how u get u=2tan(y) and du=2sec^2 ydy

OpenStudy (anonymous):

@hartnn could u continue @hba style?

OpenStudy (anonymous):

@hartnn i dont really get ur working

hartnn (hartnn):

his work is not exactly correct..

hartnn (hartnn):

ok, \(\int\limits_{}^{}\frac{ 1 }{( u+2)\sqrt{u^2+4 } }du\) u substitute here u=2tan y du=... ?

OpenStudy (shubhamsrg):

whats the problem with my solution? :D

OpenStudy (anonymous):

@hartnn ouh really?that"s y i stuck here \[\int\limits_{}^{} \frac{ du }{ u \sqrt{u^2+4} }\]

hartnn (hartnn):

oh, sorry thats is correct

hartnn (hartnn):

still u put, u=2tan y, what u get ?

hartnn (hartnn):

du=....?

OpenStudy (kainui):

@shubhamsrg I mean, your way just seems like it doesn't really do anything. It's still a trig substitution after you get it to about here: \[-\int\limits_{}^{}\frac{ dt }{ \sqrt{1+4t^2} }\]

OpenStudy (anonymous):

@shubhamsrg could you please type your work properly.,i could get it

hartnn (hartnn):

du = 2 sec^2 y dy \(\large \int \frac{2sec^2 ydy}{2 tan y.4.(secy)}=.....?\)

hartnn (hartnn):

i think there will be 2 sec y instead of 4 sec y...

OpenStudy (anonymous):

@shubhamsrg could not get it

OpenStudy (shubhamsrg):

after taking all constants aside, we get main part to integrate is 1/sqrt[ (1/2)^2 + t^2 ] which is a standard case @Kainui

hartnn (hartnn):

u got this after substitution ? \(\large \int \frac{2sec^2 ydy}{2 tan y.2.(secy)}=.....?\) ?

OpenStudy (anonymous):

@hartnn where u get u=2tan y?i cant figure it out?

hartnn (hartnn):

there was u in the denominator....

OpenStudy (kainui):

I guess I just don't see going through all that substitution as being easier, even though the result looks prettier.

OpenStudy (shubhamsrg):

@gee